OpenStudy (anonymous):

Suppose that the function f satisfies f'(x)= 3x^2-sinx(pi)x. Then the slope of the line tangent to the graph f at the point x=2 is A) 12 B)8-(1/pi) C) 7 D) 12-pi E)24

OpenStudy (anonymous):

\[f'(x)= 3x ^{2}-\sin pix\]

OpenStudy (anonymous):

\[f'(x)= 3x ^{2}-\sin (\pi x)\]This?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

Can you find the derivative?

OpenStudy (anonymous):

would it be 6x-cos(pi)x ?

OpenStudy (anonymous):

I am not sure if I have to find the derivative since its already giving me the derivative

OpenStudy (anonymous):

Oh, you're right. So just plug in \(x=2\).

OpenStudy (anonymous):

i tried that but it didnt gave me one of the answers :/

OpenStudy (anonymous):

It's one of the answers

OpenStudy (anonymous):

\[ 3(2)^2-\sin(\pi(2)=3\times 4-0=12 \]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ooh thank you:)